OpenStudy (anonymous):

f(x)=4x+7 and g(x)=3x-5. find (g o f)(-4)

OpenStudy (anonymous):

3(4x+7)-5 12x+21-5 12(-4)+21-5 -32?

OpenStudy (amistre64):

that would be a suitable process yes; or just find f(-4) and use that value in g(x)

hartnn (hartnn):

-32 correct! :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

try the alternate way shown by amistre to verify your answer

hartnn (hartnn):

find f(-4) = 4(-4)+7 = ... ?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

4(-4)+7+-9

OpenStudy (anonymous):

g(-9)=3(-9)-5=-32 :)

hartnn (hartnn):

-9 is correct, good , both of you now find g(-9) = 3(-9)-5 =... ?

OpenStudy (anonymous):

3(-9)-5=-32

hartnn (hartnn):

yup, in both ways, u get -32 which is correct :)